find n th differential coefficient of x 4 / (x-1) (x-2)

find nth differential coefficient of  x/ (x-1) (x-2)
 

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
f(x) = 1/(x^2 -4) 
1/x^2 - 4 = (x^2-4)^-1
diffrenciating using chin rule 
d/dx (f(g(x)) = f'(g(x))g'(x) 
f(x) = x^-1 
g(x) = x^2 -4 
Hope this will help 
Good Luck